OpenStudy (anonymous):

100 = 55t, solve for t. (explain the answer please)

OpenStudy (ray10):

if \[100=55t\] then: \[\frac{ 100 }{ 55} =t\]

OpenStudy (petiteme):

divide both sides by 55 :) So t=100/55

OpenStudy (ray10):

\[\frac{ 100 }{ 55 }\] simplifies down to \[\frac{ 20 }{ 11 }\]

OpenStudy (ray10):

you should keep it as a fraction because it is in its exact form :)

OpenStudy (ray10):

\[t=\frac{ 20 }{ 11 }\]

OpenStudy (anonymous):

Thankyou to everyone who helped!

OpenStudy (ray10):

anytime! :)

OpenStudy (petiteme):

no prob :) anytime!:)